OpenStudy (anonymous):

If you spend $9.74 in how many ways can you receive change for a ten-dollar bill

OpenStudy (solomonzelman):

well, the change you get is 26 cents. So, way 1. `1 quarter` + ` penny` way 2. `2 dimes` + `1 nickle` + ` penny` way 3. `1 dime` + `3 nickles` + ` penny` way 4. `5 nickles` + ` penny`

OpenStudy (solomonzelman):

So I think there are 4 ways.

OpenStudy (anonymous):

no sir there is more :) theres 26 pennies also

OpenStudy (anonymous):

i think that is it i cant think of anymore

OpenStudy (anonymous):

wait 16 pennies 1 nickel 16 pennies 1 dime

OpenStudy (anonymous):

2nickles 16 pennies

OpenStudy (anonymous):

4 nickles 6 pennies

OpenStudy (solomonzelman):

my bad So, way 1. `1 quarter` + `penny` way 2. ` 2 dimes` + `1 nickle` + `penny` way 3. ` 2 dimes` + `6 pennies` way 4. `1 dime` + `3 nickles` + `penny` way 5. `1 dime` + `2 nickles` + `6 pennies` way 6. `1 dime` + `1 nickles` + `11 pennies` way 7. `1 dime` + `16 pennies` way 8. `5 nickles` + `1 penny` way 9. `4 nickles` + `6 pennies` way 10. `3 nickles` + `11 pennies` way 11. `2 nickles` + `16 pennies` way 12. `1 nickles` + `21 pennies` way 12. `26 pennies`

OpenStudy (solomonzelman):

I am reducing, first getting all patterns with quarters, then all patterns with dimes, then with nickles and then till I get the way 13, with 26 pennies.

OpenStudy (solomonzelman):

(The last one is way 13, not way 12)

jimthompson5910 (jim_thompson5910):

SolomonZelman beat me to it, but I might as well post what I wrote $10 - $9.74 = $0.26 = 26 cents There are 13 ways to make 26 cents 1 penny + 1 quarter = 1*1 + 1*25 = 26 cents (2 coins) 1 penny + 1 nickel + 2 dimes = 1 + 5 + 2*10 = 26 cents (4 coins) 1 penny + 3 nickels + 1 dime = 1*1 + 3*5 + 1*10 = 26 cents (4 coins) 1 penny + 5 nickels = 1*1 + 5*5 = 26 cents (6 coins) 6 pennies + 2 dimes = 6*1 + 2*10 = 26 cents (8 coins) 6 pennies + 2 nickels + 1 dime = 6*1 + 2*5 + 1*10 = 26 cents (9 coins) 6 pennies + 4 nickels = 6*1 + 4*5 = 26 cents (10 coins) 11 pennies + 1 nickel + 1 dime = 11*1 + 1*5 + 1*10 = 26 cents (13 coins) 11 pennies + 3 nickels = 11*1 + 3*5 = 26 cents (14 coins) 16 pennies + 1 dime = 16*1 + 1*10 = 26 cents (17 coins) 16 pennies + 2 nickels = 16*1 + 2*5 = 26 cents (18 coins) 21 pennies + 1 nickel = 21*1 + 1*5 = 26 cents (22 coins) 26 pennies = 26*1 = 26 cents (26 coins)

OpenStudy (solomonzelman):

reducing the other direction :) I like this approach :)

OpenStudy (anonymous):

Lol thanks you guys. I appreciate it.